Solve the equations by substitution.
5x - 6 = y
2x - 3y = 4
x = ?
y = ?

Answer by checkley77(12844)

5x - 6 = y
2x - 3y = 4 replace y with (5x-6) in this equation & solve for x
2x-3(5x-6)=4
2x-15x+18=4
-13x=4-18
-13x=-14
x=-14/-13
x=14/13 ans.
5*14/13-6=y
70/13-6=y
70/13-78/13=y
y=-8/13 ans.
Proof:
2*14/13-3*-8/13=4
28/13+24/13=4
52/13=4
4=4